how can $42000 be invested, part at 9% annual simple interest and the remainder at 5% annual simple interest, so that the interest earned by the two account will be equal?
Answer by mananth(16946)

let 9% investment be x
interest = 9%x
Balance investment = 42000-x
interest = 5%(42000-x)
9%x=5%(42000-x)
9x=5(42000-x)
9x=210000-5x
14x= 210000
x= 15,000
$15,0000 at 9%
and balance at 5%